# PyRadiomics-cuda: a GPU-accelerated 3D features extraction from medical images within PyRadiomics

### v1

@@ Abstract
PyRadiomics-cuda is a GPU-accelerated extension of the PyRadiomics library, designed to address the computational challenges of extracting three-dimensional shape features from medical images. By offloading key geometric computations to GPU hardware it dramatically reduces processing times for large volumetric datasets. The system maintains full compatibility with the original PyRadiomics API, enabling seamless integration into existing AI workflows without code modifications. This transparent acceleration facilitates efficient, scalable radiomics analysis, supporting rapid feature extraction essential for high-throughput AI pipeline. Tests performed on a typical computational cluster, budget and home devices prove usefulness in all scenarios. PyRadiomics-cuda is implemented in Python and C/CUDA and is freely available under the BSD license at https://github.com/mis-wut/pyradiomics-CUDA Additionally PyRadiomics-cuda test suite is available at https://github.com/mis-wut/pyradiomics-cuda-data-gen. It provides detailed handbook and sample scripts suited for different kinds of workflows plus detailed installation instructions. The dataset used for testing is available at Kaggle https://www.kaggle.com/datasets/sabahesaraki/kidney-tumor-segmentation-challengekits-19
1 Introduction
@@
In AI era, radiomics remains an evolving discipline that seeks to extract quantitative features from standard medical images, transforming visual radiological data into high-dimensional, mineable information. This approach supports the development of predictive and prognostic models for cancer characterization, treatment response evaluation, and survival analysis. Among the most influential tools in the radiomics domain is PyRadiomics, a widely used, open-source Python library [8]. PyRadiomics provides a framework for the standardized extraction of engineered features from medical imaging data such as CT, MRI, and PET scans. These features include first-order statistics, shape descriptors, and a broad range of texture metrics derived from matrices such as GLCM, GLRLM, and GLSZM [^1] Gray Level Co-occurrence Matrix, Gray Level Run Length Matrix, Gray Level Size Zone Matrix. The library’s modular architecture, support for both 2D and 3D images, and seamless integration with platforms like 3D Slicer [1] and SimpleITK [5] have contributed to its rapid adoption in academic research and clinical trials. Since its original publication, PyRadiomics has become a cornerstone in the radiomics ecosystem, cited in over 1700 scientific publications across a wide range of applications, from lung and head-and-neck cancer studies to brain tumor segmentation and radiogenomic analysis.
Despite its strengths, one of the computational bottlenecks in PyRadiomics is the extraction of shape-based features, which rely on resource-intensive 3D geometric operations such as mesh generation and pairwise distance calculations in planes and overall shape diameters These operations are especially demanding when applied to large image volumes or high-resolution scans, often resulting in long processing times and reduced scalability which radically slows down the radiomics workflow. For instance, the computation of shape features like mesh volume, surface area, and maximum diameters can involve algorithms with theoretical complexities of $O(n^{3})$ or higher, where $n$ is the number of voxels in the Region Of Interest (ROI). This computational overhead can be a significant limitation in high-throughput settings, such as clinical trials or large-scale imaging studies, where rapid feature extraction is essential for timely analysis and decision-making.
To overcome these limitations, we present PyRadiomics-cuda, a GPU-accelerated extension of PyRadiomics that significantly improves the performance of shape feature computation by offloading key components, such as the Marching Cubes algorithm and diameter estimation, to a CUDA-enabled device. This enhancement retains full compatibility with the original PyRadiomics API, allowing seamless integration into existing pipelines while delivering substantial speedups and enabling more efficient high-throughput radiomics studies in time-sensitive settings.
1.1 Related works
@@
Several research groups have explored CUDA-enabled implementations to accelerate radiomics workflows, targeting the bottlenecks inherent in shape and texture feature extraction. One notable contribution is a CUDA-based radiomics pipeline for unsupervised medical image analysis, demonstrating performance gains over CPU-based counterparts [7]. Their approach emphasized parallel computation of texture features using CUDA kernels for high-throughput voxel-wise processing. Additionally, there have been experimental forks of PyRadiomics incorporating GPU acceleration through PyCUDA or CuPy, particularly in research settings where feature extraction needed to scale across thousands of 3D scans. However, these implementations focused only on first-order feature classes (e.g. Grey Level Matrix) and lacked full compatibility with the original PyRadiomics architecture requiring significant restructuring of input pipelines, limiting their general usability.
@@
Another significant tool in this field is cuRadiomics [3], an open-source library designed to replicate and extend PyRadiomics functionality on CUDA-enabled hardware. Unlike earlier GPU adaptations, cuRadiomics offers a comprehensive reimplementation of the radiomic feature extraction pipeline, leveraging NVIDIA’s CUDA architecture to accelerate a wide spectrum of feature classes, including first-order statistics, shape descriptors, and texture features such as GLCM and GLRLM. It incorporates memory-optimized GPU kernels, stream-based asynchronous execution, and a modular C++/CUDA architecture to support rapid feature extraction. In benchmarking studies, cuRadiomics has demonstrated speedups of 10× to 100× over traditional CPU-based methods. However, this performance gain comes with the trade-off of API divergence, as cuRadiomics does not maintain backward compatibility with PyRadiomics, often requiring tailored input formats and custom integration pipelines. It is also exclusively dedicated to the extraction of first-order statistical features (entropy, energy, etc.) and gray-level matrix features (contrast, entropy, and correlation of the matrix) and does not compute any geometric properties of the input images. Furthermore, cuRadiomics lacks meaningful integration with Python; it is merely an experimental implementation that could potentially be adapted for further use since running and configuring it is non-trivial. It also must be noted that its code repository has not been updated since 2021.
Unlike these prior efforts, PyRadiomics-cuda introduces a tightly integrated CUDA backend that maintains backward compatibility with the standard PyRadiomics API, automatically detecting and leveraging GPU resources for shape-based features without disrupting existing codebases. This design facilitates seamless adaptation in computational processes, which can be applied to large volume data sets for example in artificial intelligence learning processes. Comparing to cuRadiomics, our philosophy differs: cuRadiomics aims for full-GPU reimplementation and maximal speed only in first-order features, whereas PyRadiomics-cuda emphasizes transparent acceleration and modular integration within the original PyRadiomics framework for the 3D shape features.
2 Architecture of the system
@@
PyRadiomics-cuda is a seamlessly integrated high-performance extension of the PyRadiomics library, built to accelerate the extraction of 3D shape features using GPU computation via CUDA. It addresses the critical computational bottleneck encountered when processing large volumetric medical images, particularly in shape-based descriptors which are expensive to compute sequentially. Currently, PyRadiomics-cuda focuses exclusively on accelerating the Shape feature class within PyRadiomics, which includes metrics such as: mesh volume, surface area, maximum 3D diameter and planar diameters (XY, YZ, XZ planes).
The process of finding the above volumetric features is based on two steps. In the first one, a parallel threads in a novel highly optimized kernel create a mesh of triangles reproducing the 3d shape of the ROI in the image. Resulting vertices are passed to PyRadiomics transparently to compute all the required features using its built-in functions. However, since diameter calculation for 3D shapes is inherently complex (see Table 2), PyRadiomics-cuda also offers a parallel method for this utilizing several levels of complex parallel data gathering operations to assure the best possible performance from GPU device.
The resulting GPU kernels were optimized for performance and memory usage, leveraging shared memory, coalesced memory accesses to minimize latency and various parallel programming techniques. Overall several combinations of implementation improvements had been tested of which the most effective were added to the final version of the tool like proper organization of global memory using Structure of Arrays (SoA) and reduction of global and block-level atomic operations [9].
2.1 PyRadiomics integration
@@
The process is managed by the convenient setuptools configuration of the Python package, which has been extended to handle the CUDA compilation toolchain. During installation, the build system automatically detects the presence of the NVIDIA CUDA Compiler (nvcc). If found, it compiles the C/CUDA source files of our extension. Before the compilation, within the C extension code, a single line calling the original CPU-based shape computation function is replaced with a call to the custom dispatcher function. Later, at runtime, this dispatcher determines the execution path: it first queries for a CUDA-capable device, if a suitable GPU is available and the driver initializes successfully, the computation is offloaded to the optimized CUDA kernels. If no GPU is found or an error occurs, the dispatcher gracefully falls back to the original CPU implementation, ensuring that the library remains functional on any system. This approach allows users to benefit from GPU acceleration without any changes to their existing code. The usage is therefore the same as in the other PyRadiomics applications and may be as simple as:
from radiomics import featureextractor
ext = featureextractor.RadiomicsFeatureExtractor()
res = ext.execute(’scan.nii.gz’, ’mask.nii.gz’) #input
print(res[’MeshVolume’], res[’SurfaceArea’])
3 Efficiency experiments and runtime results
3.1 Test framework
To validate the numerical correctness and rigorously benchmark the performance of various optimization strategies, a standalone test framework was developed in C and CUDA. It allows for the direct comparison of multiple algorithm implementations, including the original CPU baseline and numerous versions of the GPU kernels, each employing different optimization techniques. A key feature of the test harness is its capacity for granular time measurement, profiling not only the total execution time but also the time consumed by distinct computational stages, such as the Marching Cubes mesh generation and the subsequent diameter calculations, independently.
In order to ensure data integrity between the Python environment and the C test framework, a Python script intercepts the data within the PyRadiomics feature extractor just before it is passed to the C computation function. This script serializes the untouched NumPy arrays (image and mask) to disk as .npy files. The C test framework then utilizes NumPy C API headers to parse these files, guaranteeing that the exact binary data is loaded for testing. The parsed data is passed directly to the C function calls, and execution times are precisely measured using high-resolution system timers, facilitating an accurate and reliable comparison of all implemented algorithms.
All the tests included the following pipeline: reading the input images, applying the segmentation masks computing the radiometric features including volume, surface area and diameters.
3.2 Hardware used and testing data set
There were three different machines used: (1) a professional modern GPU cluster built for AI learning, (2) an average desktop computer and (3) an old server machine equipped with a budget GPU designed to enable AI in refurbished hardware. These configurations are displayed in the Table 1. As an input data set we used Kidney Tumor Segmentation Challenge 19 (KITS19) [2]. The complete set contains more than 200 cases. However, we selected 20 samples covering exemplification of the possible input images sizes from 50kB to 9MB and producing from $2700$ to $236588$ 3-dimensional feature vertices.
@@ 3.2 Hardware used and testing data set
3.3 Discussion of the results
In Fig. 1 we can see processing times of the KITS19 data set using PyRadiomics and PyRadiomics-cuda on various hardware configurations with a workflow including creation of a 3D shape from given images and ROI and then calculating volumetric features, diameters and area. The time is presented on log-log scale due to the large differences between the processing times for various datasets from less than 1 millisecond to about 2 minutes: 59 milliseconds on PyRadiomics-cuda on H100 GPU compared to 121 seconds for PyRadiomics on Intel Xeon on 236588 vertices and the same input data.
Figure 1: Time efficiency of the KITS dataset feature extraction using PyRadiomics and PyRadiomics-cuda on the various machines. (See Supplementary Information for exact data values of this plot.)
A detailed analysis of the workflow in PyRadiomics indicated that most of the time after reading input file is spent in diameter calculation, from 95.7% for small images to 99.9% for large ones (see Table 2). This is expected due to diameter calculation’s theoretical algorithmic complexity of $O(n^{4})$. The Marching Cubes algorithm, with a complexity of $O(n^{3})$, was significantly faster but still took non-negligible time for large ROIs. Reducing these times was the primary aim of the optimization. Since PyRadiomics is not able to utilize multiple CPU cores, switching to better CPUs lead to only limited speedup, which could be attributed to clock speed and architecture improvements. This can be easily observed in Fig. 2, where switching between CPU configurations 1, 2 and 3 did not reduce the processing time more than 3 times. Using parallel processing on GPUs led to much higher speedup of Marching Cubes and diameter calculation algorithms, which can be attributed to use of massively parallel threads. Even on a budget GPU, PyRadiomics-cuda achieved a speed-up from 8 to 24 times in 3D features extraction. For modern GPUs and more expensive hardware, the processing time was reduced by more than 50 and even 2000 times when compared to our Intel Xeon processor.
In Table 2 we can also observe that PyRadiomics spends a long time loading data files. This cost comes not only from disk operations but also from various other operations like cleaning, normalization and conversion to proper memory layout. Improving this step still remains an open challenge which could be done on GPU as well.
For complete workflows including data loading and preprocessing utilization of PyRadiomics-cuda resulted in 8 times speed-up for bigger data files with more than 200k vertices in 3D space. Since this is not uncommon case for bigger ROIs, we predict huge savings in processing costs for machine learning tasks on clusters.
Figure 2: Available speedup of 3D features processing when using PyRadiomics-cuda on various GPU devices compared to the original implementation working on CPU processors. Intel Xeon with PyRadiomics taken as a base reference time. (See Supplementary Information for exact data values of this plot.)
@@ 3.3 Discussion of the results
Table 2: The breakdown of the processing time in milliseconds into individual steps for PyRadiomics on CPU processor and PyRadiomics-cuda on GPU processor, both from conf. 2 in Table 1. M.Cubes – Marching Cubes algorithm time, Diameters – 3D diameters calculation time, D. transfer – time needed to copy data from CPU to GPU memory. Representative sample of KITS’19 data set [2]. (Complete table data available in Supplementary Information.)
@@
case image vertices in File [ms] PyRadiomics time [ms] PyRadiomics-cuda time [ms] Speedup [x1]
no. size 3D space reading M.Cubes Diameters Total D. transfer M.Cubes Diameters Total comp. overall
00000-1 231x104x264 124406 2346 20.702 9516.538 9537.24 8.045 7.179 514.797 530.021 17.994 4.131
00000-2 28x30x59 6132 2350 0.362 25.256 25.618 0.326 0.184 2.418 2.928 8.75 1.01
00001-1 322x126x219 236588 2494 29.451 34210.327 34239.778 9.73 11.006 1855.825 1876.561 18.246 8.404
00001-2 51x62x135 8928 2521 2.274 51.383 53.657 0.683 0.554 3.432 4.668 11.494 1.019
00002-1 230x109x163 83098 1032 13.385 4256.214 4269.599 5.065 4.811 231.794 241.671 17.667 4.162
00002-2 50x45x44 9206 1024 0.577 56.903 57.48 0.453 0.29 3.94 4.683 12.274 1.051
00003-1 237x122x135 77560 1105 12.69 3730.998 3743.688 4.824 4.622 204.071 213.517 17.533 3.677

### v2

# PyRadiomics-cuda: 3D features extraction from medical images for HPC using GPU acceleration

@@ Abstract
PyRadiomics-cuda is a GPU-accelerated extension of the PyRadiomics library, designed to address the computational challenges of extracting three-dimensional shape features from medical images. By offloading key geometric computations to GPU hardware it dramatically reduces processing times for large volumetric datasets. The system maintains full compatibility with the original PyRadiomics API, enabling seamless integration into existing AI workflows without code modifications. This transparent acceleration facilitates efficient, scalable radiomics analysis, supporting rapid feature extraction essential for high-throughput AI pipeline. Tests performed on a typical computational cluster, budget and home devices prove usefulness in all scenarios.
1 Introduction and related works
@@
Radiomics extracts numeric features from medical images, turning images into analyzable data. These features help build predictive models for cancer diagnosis, treatment response, and survival. PyRadiomics is a widely used open-source Python library for radiomics [van2017computational]. It standardizes feature extraction from CT, MRI, and PET scans. They include first-order stats, shape descriptors, and texture metrics (GLCM, GLRLM, GLSZM [^1] Gray Level Co-occurrence Matrix, Gray Level Run Length Matrix, Gray Level Size Zone Matrix). Its modular design, 2D/3D support, and integration with tools like 3D Slicer [slicer] and SimpleITK [simpleitk] drove wide academic and clinical adoption. Since release, PyRadiomics has been cited over 1700 times across many medical imaging studies.
However, shape-based features are computationally heavy, requiring 3D operations like mesh generation and pairwise distance calculations. On large or high-resolution scans these steps are slow and limit scalability. For example, mesh volume and surface area can be $O(n^{3})$ and diameter calculations can be $O(m^{2})$, with $n$ and $m$ the image linear resolution and ROI voxel count. We found PyRadiomics can spend up to 99.9% of time on diameter calculations for large ROIs (see Table 2), a major bottleneck for large-scale AI workflows.
@@
Several research groups have explored CUDA-enabled implementations to accelerate radiomics workflows, targeting the bottlenecks in shape and texture feature extraction. Rundo et al. presented a CUDA-based radiomics pipeline for unsupervised medical image analysis, demonstrating performance gains over CPU-based counterparts. Their approach emphasized parallel computation of texture features using CUDA kernels for high-throughput voxel-wise processing [rundo_cuda-powered_2021].
There have been experimental forks of PyRadiomics incorporating GPU acceleration through PyCUDA or CuPy, particularly in research settings where feature extraction needed to scale across thousands of 3D scans. However, these implementations focused only on first-order feature classes (e.g., Grey Level Matrix), lacked full compatibility with the original PyRadiomics architecture (requiring significant restructuring of input pipelines), limited general usability, and are no longer maintained.
@@
Another significant tool in this field is cuRadiomics [jiao_curadiomics_2020], an open-source library designed to replicate and extend PyRadiomics functionality on CUDA-enabled hardware. Unlike earlier GPU adaptations, cuRadiomics offers a comprehensive reimplementation of the radiomic feature extraction pipeline, leveraging NVIDIA’s CUDA architecture to accelerate a wide spectrum of feature classes, including first-order statistics, shape descriptors, and texture features such as GLCM and GLRLM. It incorporates memory-optimized GPU kernels, stream-based asynchronous execution, and a modular C++/CUDA architecture to support rapid feature extraction. In benchmarking studies, cuRadiomics has demonstrated speedups of 10× to 100× over traditional CPU-based methods. However, this performance gain comes with the trade-off of API divergence, as cuRadiomics does not maintain backward compatibility with PyRadiomics, often requiring tailored input formats and custom integration pipelines. It is also exclusively dedicated to the extraction of first-order statistical features (entropy, energy, etc.) and gray-level matrix features (contrast, entropy, and correlation of the matrix) and does not compute any geometric properties of the input images. Furthermore, cuRadiomics lacks meaningful integration with Python; it is merely an experimental implementation that could potentially be adapted for further use since running and configuring it is non-trivial. Its code repository has not been updated since 2021 and can hardly run with new CUDA and GPUs.
Unlike these prior efforts, PyRadiomics-cuda:
  • adds a CUDA backend that preserves PyRadiomics compatibility;
  • auto-detects GPUs and falls back to CPU if no GPU is found;
  • transparently accelerates 3D shape features.
Obviously 3D features extraction is not the only bottleneck in radiomics workflows, with data reading and data transfer between CPU and GPU also contributing significantly to processing time. In our research we analyzed the time spent on different steps of the workflow and justify PyRadiomics-cuda utilization in various configurations.
2 PyRadiomics-cuda design
@@
From computational point of view, as it is visible in Table 2 in case of large 3D volumes, PyRadiomics processing speed is limited by 3D features extraction time. Therefore, PyRadiomics-cuda focuses exclusively on accelerating the Shape feature class within PyRadiomics, which includes metrics such as: mesh volume, surface area, maximum 3D diameter and planar diameters (XY, YZ, XZ planes).
The process of finding the above volumetric features is based on two steps. Firstly it creates a mesh of triangles reproducing the 3D shape of the ROI in the image and on the fly calculates the volume and surface area of the shape. In the second step it walks again the triangles in order to find the maximum 3D diameter and planar diameters.
Marching cubes fused parallel kernels.
PyRadiomics uses a brute-force version of the Marching Cubes algorithm [lorensen1987marching] for 3D mesh generation. It walks through all voxels in the 3D image, checking each voxel’s eight corner values against a specified isovalue (typically the segmentation threshold). Number of voxels containing this isosurface is usually much smaller than the total number of voxels in the image, so this approach is efficient for smaller cases. Our parallel version of this process in general follows the same logic, but each voxel is processed in a separate thread. If a thread detects that the isosurface passes through its voxel, it computes the corresponding triangle and adds it to the global list of vertices and also updates the total volume and surface area. Obviously this process requires synchronization between threads to ensure that the global data structures are updated correctly and here several possible approaches were tested (see Section 3) in order to find the best one for different GPU architectures.
Parallel diameter calculation.
Calculating the maximum 3D diameter of a shape is computationally intensive, as it finds the pair of vertices that are the farthest apart.
The original PyRadiomics implementation suffers from inefficiency, especially for large meshes with many vertices. To optimize this process for GPU execution, we again implemented a parallel algorithm that divides the workload among multiple threads. Each thread computes the distances between a subset of vertex pairs, storing the maximum distance found in an accumulator. After all threads complete their calculations, a reduction operation is performed to find the overall maximum distance from the local maxima.
There are several ways to implement parallel accumulation process and we tested various strategies to find the best one for different GPU architectures (see Section 3).
The resulting GPU kernels were optimized for performance and memory usage, leveraging shared memory, coalesced memory accesses to minimize latency and various parallel programming techniques [cuda_handbook].
PyRadiomics integration.
@@
The process is managed by the convenient setuptools configuration of the Python package, which has been extended to handle the CUDA compilation toolchain. During installation, the build system automatically detects the presence of the NVIDIA CUDA Compiler (nvcc). If found, it compiles the C/CUDA source files of our extension. Before the compilation, within the C extension code, a single line calling the original CPU-based shape computation function is replaced with a call to the custom dispatcher function. Later, at runtime, this dispatcher determines the execution path: it first queries for a CUDA-capable device, if a suitable GPU is available and the driver initializes successfully, the computation is offloaded to the optimized CUDA kernels. If no GPU is found or an error occurs, the dispatcher gracefully falls back to the original CPU implementation, ensuring that the library remains functional on any system. This approach allows users to benefit from GPU acceleration without any changes to their existing code. The usage is therefore the same as in the other PyRadiomics applications and is as simple as:
[…]
,fontsize=,numbers=left,numbersep=5pt,xleftmargin=0.3cm,,
from radiomics import featureextractorext = featureextractor.RadiomicsFeatureExtractor()res = ext.execute(’scan.nii.gz’, ’mask.nii.gz’) #inputprint(res[’MeshVolume’], res[’SurfaceArea’])
3 Optimization research and results
Hardware used, testing data set and test framework.
There were three different machines used: (1) a professional modern GPU cluster built for AI, (2) an average desktop computer and (3) an old server machine equipped with a budget GPU designed to enable AI in refurbished hardware (specs listed in Table 1). As an input data set we used public data from Kidney Tumor Segmentation Challenge 19 (KITS19) [heller2020kits19, KidneyTumorSegmentation]. The complete set contains more than 200 cases, but we selected 20 samples covering exemplification of the possible input images sizes from 50kB to 9MB and producing from $2700$ to $236588$ 3D feature vertices.
@@ 3 Optimization research and results
Micro-benchmarking the optimizations.
In order to find the PyRadiomics performance and the best possible combination of techniques that would yield the highest improvement for various GPU hardware and also rigorously validate the correctness of the results, we created a micro-benchmarking tool [MiswutPyradiomicscudadatagen2026]. It directly compares multiple algorithm implementations, including the original CPU baseline and numerous versions of the GPU kernels, each employing different optimization techniques. All benchmarks use the same pipeline: reading the input images, applying the segmentation masks, transferring data to GPU (if needed), computing the features: volume, surface area and diameters.
Analysis of PyRadiomics in processing 3D shape features revealed that the main bottleneck appears in diameter calculation (see Table 2), therefore we focused our optimization research on this part of the algorithm.
@@
Table 2: The breakdown of the processing time in milliseconds into individual steps for PyRadiomics on CPU processor and PyRadiomics-cuda on GPU processor, both from desktop conf. 2 in Table 1. M.C. – Marching Cubes algorithm time, Diam. – 3D diameters calculation time, D. tran. – time needed to copy data from CPU to GPU memory, Comp. – speedup in pure computations, Overall – speedup including input file reading. Representative sample of KITS’19 data set [heller2020kits19].
@@
case image vertices File PyRadiomics time PyRadiomics-cuda time Speedup
no. size in 3D reading M.C. Diam. Total D. tran. M.C. Diam. Total Comp. Overall
space [ms] [ms] [ms] [x1]
00000-1 231x104x264 124406 2346 20.7 9516.5 9537.2 8.0 7.2 514.8 530.0 18.0 4.1
00000-2 28x30x59 6132 2350 0.4 25.3 25.6 0.3 0.2 2.4 2.9 8.8 1.0
00001-1 322x126x219 236588 2494 29.5 34210.3 34239.8 9.7 11.0 1855.8 1876.6 18.2 8.4
00001-2 51x62x135 8928 2521 2.3 51.4 53.7 0.7 0.6 3.4 4.7 11.5 1.0
00002-1 230x109x163 83098 1032 13.4 4256.2 4269.6 5.1 4.8 231.8 241.7 17.7 4.2
00002-2 50x45x44 9206 1024 0.6 56.9 57.5 0.5 0.3 3.9 4.7 12.3 1.1
00003-1 237x122x135 77560 1105 12.7 3731.0 3743.7 4.8 4.6 204.1 213.5 17.5 3.7
00003-2 39x35x31 4568 1097 0.2 14.7 14.9 0.3 0.2 1.6 2.1 7.1 1.0
00004-1 254x70x36 31838 254 2.5 677.2 679.7 0.8 1.1 37.8 39.7 17.1 3.2
00004-2 35x37x10 2742 255 0.1 5.7 5.8 0.3 0.1 1.1 1.4 4.0 1.0
00005-1 167x94x285 126446 3150 15.0 9780.9 9795.9 5.6 5.6 531.5 542.7 18.1 3.5
00005-2 51x53x121 22024 3203 1.9 305.6 307.4 0.6 0.7 18.0 19.3 15.9 1.1
00006-1 308x102x36 65436 710 4.4 2828.1 2832.5 1.1 2.0 153.7 156.8 18.1 4.1
00006-2 41x43x13 3676 712 0.1 10.0 10.1 0.3 0.2 1.1 1.6 6.5 1.0
00007-1 265x101x39 49912 255 4.1 1634.9 1638.9 1.0 1.7 90.1 92.8 17.7 5.4
00007-2 39x43x12 3498 250 0.1 9.3 9.5 0.3 0.1 1.2 1.6 6.0 1.0
00008-1 288x177x54 57362 967 9.3 2089.4 2098.6 3.3 3.1 113.7 120.1 17.5 2.8
00008-2 127x154x41 47484 972 3.2 1436.9 1440.2 0.8 1.4 78.7 80.9 17.8 2.3
00009-1 241x95x47 37576 337 3.8 916.2 920.1 1.1 1.5 50.5 53.0 17.4 3.2
00009-2 39x33x11 2700 340 0.1 5.7 5.8 0.3 0.1 1.1 1.5 3.9 1.0
The initial implementation serving as a baseline (1) included basic techniques and equal threads load-balancing. The second optimization (2) added block-based atomic reductions. The third step improved theoretical memory access time (3) by using 2D structures loaded in the shared memory. Since this optimization was not effective on older hardware with less shared memory per block, we introduced local thread accumulators (4) to reduce the number of atomic operations. Finally, we simplified the memory access patterns (5) to use just 1D arrays instead of 2D structures, which reduced the number of calculations needed for indexing. We found out that the conflicts between threads are rare, so the overhead of the simplest possible atomic operations is negligible.
In Fig. 1 we can see the comparison of the processing times for all the above-mentioned optimizations on tested GPU devices. We observed that in case of an older T4 device atomic operations are not as effective and block-based reduction is the best strategy. Modern H100 offers fast atomic operations but needs more attention when accessing global memory. Finally, RTX4070 was the most effective with local thread accumulators. Simplified memory accesses by using 1D arrays did not improve performance significantly on any device when compared to other ideas, so it was not included in the final implementation.
Figure 1: Comparison of optimization techniques applied to the GPU kernels for shape feature extraction across three GPU models. X-axis: sum of processing time of all input files in milliseconds on a logarithmic scale.
Discussion of the results.
In Fig. 2 we can see processing times of the KITS19 data set using PyRadiomics and PyRadiomics-cuda on various hardware configurations within a computational workflow including creation of a 3D shape from given images and ROI and then calculating volumetric features, diameters and area. The time is presented on log-log scale due to the large differences between the processing times for various datasets from 59 milliseconds on PyRadiomics-cuda on H100 GPU compared to 121 seconds for PyRadiomics on Intel Xeon on 236588 vertices and the same data.
Figure 2: LEFT: Time efficiency of the feature extraction using PyRadiomics and PyRadiomics-cuda on the various machines. RIGHT: Available speedup of 3D features processing using PyRadiomics-cuda on various GPU devices compared to the original implementation working on CPU. Intel Xeon with PyRadiomics taken as a base reference time.
A detailed analysis of the workflow in PyRadiomics indicated that most of the time after reading input file is spent in diameter calculation, from 95.7% for small images to 99.9% for large ones (see Table 2). Reducing these times was the primary aim of the optimization. Since PyRadiomics is not able to utilize multiple CPU cores, switching to better CPUs lead to only limited speedup, which could be attributed to the clock speed and architecture improvements. This can be easily observed in Fig. 2, where switching between CPU configurations 1, 2 and 3 did not reduce the processing time more than 3 times. Using parallel processing on GPUs led to much higher speedup, which can be attributed to use of massively parallel threads. Even on a budget GPU, PyRadiomics-cuda achieved a speed-up from 8 to 24 times in 3D features extraction. For modern GPUs and more expensive hardware, the processing time was reduced by more than 50 and even 2000 times, compared to Intel Xeon processor.
In Table 2 we can also observe that PyRadiomics spends a long time loading data files. This cost comes not only from disk operations but also from various other operations like cleaning, normalization and conversion to proper memory layout. Improving this step still remains an open challenge which could be done on GPU as well. In all processed files, data transfer between CPU and GPU was compensated by faster processing and we did not observe any performance degradation. However, for smaller files the time spent on reading files was dominating the overall performance and no observable speedup was achieved.
For complete workflows (including data loading and preprocessing) performed on a desktop machine utilization of PyRadiomics-cuda resulted in 8 times speed-up for bigger data files with more than 200k vertices in 3D space. Direct Memory Access and fast data transfer will lead To huge savings in processing costs for machine learning tasks on clusters.